OpenStudy (anonymous):

Help please!!!!!!! The length of a text messaging conversation is normally distributed with a mean of 2 minutes and a standard deviation of .5 minutes. What is the probability that a conversation lasts longer than 3 minutes?

OpenStudy (anonymous):

i think that's the first step\[1.96*\sqrt{\frac{ .5(1-.5) }{ 3 }}\]

OpenStudy (kropot72):

First you need to find the z-score for 3 minutes: \[\large z=\frac{X-\mu}{\sigma}=\frac{3-2}{0.5}=2\] Now you can use a standard normal distribution table to find the cumulative probability for a z-score of 2, which is 0.9772. So we now know that the probability a conversation lasts 3 minutes or less is 0.9772. Therefore the probability that a conversation lasts longer than 3 minutes is given by: P(longer than 3 minutes) = 1.0000 - 0.9772 = you can calculate

OpenStudy (anonymous):

you're awesome. thanks so much :D

OpenStudy (kropot72):

You're welcome :)